An unknown gas at 75.1 ∘C and 1.00 atm has a molar mass of 30.07 g/mol. Assuming ideal behavior, what is the density of the gas?

density: ______ g/L
Answer #1

P * molar mass = density * R * T

where, P = pressure = 1.00 atm

molar mass = 30.07 g/mol

R = Gas constant

T = temperature = 75.1 +273 = 348.1 K

1.00 * 30.07 = density * 0.0821 * 348.1

30.07 = density * 28.6

density = 30.07 / 28.6 = 1.05 g/L

Therefore, the density of gas = 1.05 g/L